Sodium carbonate is a reagent that may be used to standardize acids in the same way you used 𝐾𝐻𝑃 in this experiment. In such a standardization, it was found that a 0.512 𝑔 sample of sodium carbonate required 26.30 π‘šπΏ of a sulfuric acid solution to reach the end point for the reaction. π‘π‘Ž2𝐢𝑂3 (π‘Žπ‘ž) + 𝐻2𝑆𝑂4 (π‘Žπ‘ž) β†’ 𝐻2𝑂 (𝑙) + 𝐢𝑂2 (𝑔) + π‘π‘Ž2𝑆𝑂4(π‘Žπ‘ž) What is the molarity of the 𝐻2𝑆𝑂4?Β 


1
Expert's answer
2022-05-17T13:11:03-0400

Answer:

M = 0.1825 M

Explanation:

To do this, let's write the equation again:

Naβ‚‚CO₃ + Hβ‚‚SOβ‚„ ---------> Hβ‚‚O + COβ‚‚ + Naβ‚‚SOβ‚„

As we can see, the equation is already balanced and we can also see that the mole ratio between the acid and the carbonate is 1:1, this means that the moles of the acid, would be the same moles of the carbonate, therefore, we can use the following expression:

M₁V₁ = Mβ‚‚Vβ‚‚ (1)

1: Is the carbonate

2: is the acid

To get the concentration of the acid, we need to calculate the moles of the carbonate used. This can be done using the molecular mass of the sodium carbonate, which is 105.9888 g/mol, so the moles:

n₁ = 0.512 / 105.9888 = 0.0048 moles

Now that we have the moles, we can use (1) and calculate the concentration of the acid.

We know that:

n₁ = M₁V₁ (2)

Replacing in (1) we have:

n₁ = Mβ‚‚Vβ‚‚

Mβ‚‚ = n₁ / Vβ‚‚ (3)

Now all we have to do is replace the values and solve for the concentration:

Mβ‚‚ = 0.0048 / (0.02630)

Mβ‚‚ = 0.1825 MThis is the concentration (molarity) of the Hβ‚‚SOβ‚„